OpenStudy (anonymous):

Solve the equation: 2- 2x = sqrt{4 + x} I tried solving this, and I got x = 9/4, 0. But the answer to this is only x = 0. Why?

OpenStudy (anonymous):

look at this plot

OpenStudy (anonymous):

2-2(9/4) is a negative number. The Square root side can not equal a negative number. Therefore 9/4 is extraneous

OpenStudy (anonymous):

Plug in 9/4 to verify and you'll see why that cant be an answer.....(the suspense!!!!)

OpenStudy (anonymous):

mrjoselson ruined the ending <.< haha jk jk

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Why can't it equal a negative number? I thought square roots had that ± sign in front of them..

OpenStudy (anonymous):

The function: \[f(x) = \sqrt{x}\] always returns the positive root. If it returned both the positive and negative, it wouldnt be a function anymore (it would fail the vertical line test). Thats why when we want the negative root, or to show its existance we put a minus, or plus/minus in front.

OpenStudy (anonymous):

i hope myininaya has a better explanation lol >.<

OpenStudy (anonymous):

Here is a complete solution with explanations:

OpenStudy (anonymous):

I wouldnt use the complete the square technique, thats a little too complicated for this problem: \[4x^{2}-9x = x(4x-9)=0 \rightarrow x = 0, 4x-9=0 \rightarrow x = 0, x = \frac{9}{4}\] Other than that, beautiful :) medals for everyone >.>

myininaya (myininaya):

\[2-2x=\sqrt{4+x}\] \[(2-2x)^2=(\sqrt{4+x})^2\] \[(2-2x)^2=4+x\] \[(2-2x)(2-2x)=4+x\] \[4-4x-4x+4x^2=4+x\] \[4x^2-8x+4=4+x\] \[4x^2-8x-x+4-4=0\] \[4x^2-9x=0\] \[x(4x-9)=0\] \[x=0, x=\frac{9}{4}\] Check x=0: \[2-2(0)=2=2=\sqrt{4+0}\] so x=0 works! Check \[x=\frac{9}{4}\]: \[2-2(\frac{9}{4})=2-\frac{9}{2}=\frac{8}{4}-\frac{9}{4}=\frac{-1}{4}\] \[\sqrt{4+\frac{9}{4}}=\sqrt{\frac{16}{4}+\frac{9}{4}}=\sqrt{\frac{25}{4}}=\frac{5}{2}\] so we don't have the samething on both sides so \[x=\frac{9}{4}\] is not a solution

myininaya (myininaya):

whenever you raise both sides to an even power be sure to check your answers